# Strong order-one convergence of the Euler method for random ordinary differential equations driven by semi-martingale noises

### v7

@@ Abstract.
It is well known that the Euler method for a random ordinary differential equation $\mathrm{d}X_{t}/\mathrm{d}t=f(t,X_{t},Y_{t})$ driven by a stochastic process $\{Y_{t}\}_{t}$ with $\theta$-Hölder sample paths is estimated to be of strong order $\theta$ with respect to the time step, provided $f=f(t,x,y)$ is sufficiently regular and with suitable bounds. This order is known to increase to $1$ in some special cases. Here, it is proved that, in many more typical cases, further structures on the noise can be exploited so that the strong convergence is of order 1. More precisely, we prove so for any semi-martingale noise. This includes Itô diffusion processes, point-process noises, transport-type processes with sample paths of bounded variation, and time-changed Brownian motion. The result follows from estimating the global error as an iterated integral over both large and small mesh scales, and by switching the order of integration to move the critical regularity to the large scale. The work is complemented with numerical simulations showing the optimality of the strong order 1 convergence in those cases, and with an example with fractional Brownian motion noise with Hurst parameter $0<H<1/2,$ which is not a semi-martingale and for which the order of convergence is $H+1/2$, hence lower than the attained order 1 in the semi-martingale case, but still higher than the order $H$ of convergence expected from previous works.
@@
The second author was partly supported by Laboratório de Matemática Aplicada, Instituto de Matemática, Universidade Federal do Rio de Janeiro (LabMA/IM/UFRJ), Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES), Brasil, grant 001, and Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brasília, Brasil, grant no. 408751/2023-1
@@ 1. Introduction
$\begin{cases}\displaystyle\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=f(t,X_{t},Y_{t}),&0\leq t\leq T,\\ \left.X_{t}\right|_{t=0}=X_{0},\end{cases}$ (1.1)
on a time interval $I=[0,T]$, with $T>0$, where the initial condition $X_{0}$ is a given random variable, and the noise $\{Y_{t}\}_{t\in I}$ is a given stochastic process, both defined on a sample space $\Omega$ with probability distribution $\mathbb{P}.$ The function $f=f(t,x,y)$ is such that the differential equation in Equation 1.1 can be either a scalar or a system of equations and the noise can also be either a scalar or a vector-valued process.
Random ODEs appear in applications where the parameters either vary randomly or are not fully known. As such, they have been playing an important role in modeling practical problems and in uncertainty quantification (see e.g. [3, 9, 15, 18, 27, 30, 34, 35]). They are, in some cases, milder versions of the more popular stochastic diffusion equations (SDE), but, depending on the situation, they may actually lead to more realistic models, where, for example, the white noise is replaced by a more structured noise.
@@
An approximation $\{X_{t_{j}}^{N}\}_{j}$ is said to converge to $\{X_{t}\}_{t}$ with strong order $\theta>0$ when there exists a constant $C\geq 0$ such that
@@
[…]
@@
Under certain regularity conditions on $f$ and when the noise $\{Y_{t}\}_{t\in I}$ has $\theta$-Hölder continuous sample paths, the result in [39, Theorem 3] estimates the order of convergence of the Euler scheme, in the mean square sense, to be of order $\theta$ with respect to the time step. This implies the strong convergence Equation 1.3 with the same order $\theta$. In the case of fractional Brownian motion noise with Hurst parameter $H$, [39, Theorem 2] estimates the mean square convergence to be of order $H$, for any $0<H\leq 1$
In some special cases, the order of convergence of the Euler method is known to be higher than the Hölder exponent of the noise. For example, for Itô diffusion noises, the RODE can be seen as part of a system of SDEs, with the Euler method for the RODE component coinciding with the Milstein method for the SDE, which is known to be of strong order 1 (see [39, Section 5], in particular [39, Section 5.2, Example 12 and Remark 13]).
Figure 1. (a) Estimated order of convergence of the strong error of the Euler method for the linear system Equation 1.4 as it depends on the noise, starting with all the noises combined (all) and continuing with the scalar equation, each with a different type of noise process, namely Wiener (W), Ornstein-Uhlenbeck (OU), geometric Brownian motion (gBm), a homogeneous linear Itô diffusion (hlp), compound Poisson process (cP), step point process (sP), Hawkes process (H), transport process (T), and also a fractional Brownian motion with Hurst parameter $11/2<H<$ 1 . (fBm); (b) Sample paths of all the noises used in the linear system Equation 1.4.
In many more classical examples, though, the actual order of convergence seems to be higher than what the theory predicts. Our aim here is to show that, indeed, it is possible to exploit further properties of the noise that yield a strong order 1 convergence for essentially any noise that is a semi-martingale (see Theorem 7.1 for the precise statement). This includes noises such as point processes, transport processes, Itô processes and time-changed Brownian motions, some of which are not even continuous and for which the previous theories did not apply.
This strong order 1 convergence is not only a major improvement from previous results, but it is also remarkable that it follows with a unified proof for all semi-martingale noises. It also has potential extensions to pathwise convergence for RODEs in the case of non-globally Lipschitz functions, at least when driven by finite-variation process. The result should also extend to random partial differential equations. The technique may also yield improvements in the results concerning higher order methods and or other types of noises, although this requires a more thorough investigation.
@@
Figure 3. (a) Estimated order of convergence of the strong error of the Euler method for the toggle-switch model Equation 1.6; (b) Euler approximations of the component $\{X_{t}\}_{t}$ of a sample path.
We numerically verify, in a series of examples, that the strong order 1 convergence is in fact optimal. Indeed, this work was motived by the fact that the observed convergence in these and many other examples was higher than what the theory predicted. The examples are presented in more details in the supplementary material in Section 8 and in the github repository [23] and include the following:
@@
    where $\{Y_{t}\}_{t}$ is either (i) a univariate process of one of many types of semi-martingale processes or (ii) a vector-valued process combining all these noises as independent components.
@@
    A logistic model of population dynamics with random coefficients, loosely inspired by [18, Section 15.2], where the specific growth is the sine of a geometric Brownian motion process $\{G_{t}\}_{t}$ and with an extra point-process random term $\{H_{t}\}_{t}$ representing harvest:
@@
    A toggle-switch model of gene expression (similar to [3, Section 7.8], originated from [38], see also [35]) driven by a combination of a compound Poisson point process $\{A_{t}\}_{t}$ and an Itô diffusion process $\{B_{t}\}_{t},$
@@
    for the surplus $U_{t}=X_{t}+C_{t}+O_{t}$ of an insurance company, inspired by [16] and [10], where $C_{t}=\sum_{i=1}^{N_{t}}C_{i}$ defines a compound Poisson process modeling the insurance claims, $\{R_{t}\}_{t}$ is a geometric Brownian motion process modeling the interest rate, and $\{O_{t}\}_{t}$ is an Ornstein-Uhlenbeck process, used to transform an additive SDE into a RODE.
@@ A Fisher-KPP partial differential equation
    as inspired by the works [34] and [15] (see also [14] and [25]), and where the noise $\{Y_{t}\}_{t}$ is an Ornstein-Uhlenbeck process modulated by a decaying self-exciting Hawkes point process.
@@
    where $\{B^{H}_{t}\}_{t}$ is a fractional Brownian motion (fBm) with Hurst parameter $10<H<$ 1, which is a Wiener process when $H=1/2$ and which is not a semi-martingale for $H\neq 1/2.$
Figure 4. (a) Estimated order of convergence of the strong error of the Euler method for the insurance risk model Equation 1.8; (b) Sample paths of the surplus $\{U_{t}\}_{t},$ the Ornstein-Uhlenbeck process $\{O_{t}\}_{t},$ the interest rate $\{I_{t}\}_{t},$ and the insurance claims $\{C_{t}\}_{t}.$
Figure 5. (a) Estimated order of convergence of the strong error of the Euler method for the Fisher-KPP model Equations 1.9 and 1.10; Sample paths of (b) the Ornstein-Uhlenbeck and Hawkes processes, and of (c) the noise obtained as the product of these processes.
In all but the last example, the numerically obtained convergence was found to be of strong order approximately 1, while previous estimates would either not be applicable or yield a much lower order of convergence (see Figures 1, 2, 3, 4 and 5).
@@
The starting point of the proof is to not estimate the local error and, instead, work with an explicit formula for the global error (see Lemma 5.1), similarly to the way it is done for approximations of stochastic differential equations. In this way, we obtain the estimate (see Equation 6.3)
@@
The only problematic, noise-sensitive term in Equation 1.12 is the last one. The classical analysis is to use an assumed $\theta$-Hölder regularity of the noise sample paths and estimate the local strong error as $C\Delta t_{N}^{\theta}.$ Instead, we look at the global noise error and show that the steps of the process given by $F_{t}=f(t,X_{\tau^{N}(t)}^{N},Y_{t})$ can be controlled in a suitable global way. In order to give the main idea, let us consider a scalar equation with a scalar noise and assume that the sample paths of $\{F_{t}\}_{t\in I}$ satisfy
@@
For a semi-martingale noise, the integral above can be a combination of a Lebesgue-Stieltjes integral, a stochastic integral, and an integral with respect to a jump measure, but we leave it like that just for the sake of explaining the main idea. With that, we bound the global error term using the Fubini Theorem, taking into consideration that, despite the inner integral being from $\xi$ to the “future” time $\tau^{N}(\xi)+\Delta t_{N}$, the integrand is still non-anticipative, hence
@@
In the fractional Brownian motion case, which is not a semi-martingale, the above estimate does not apply. Instead, we essentially have (see Equation 8.12 and Equation 8.14)
$F_{s}-F_{\tau}\sim\int_{\tau}^{s}(s-\tau)^{H-1/2}\;\mathrm{d}W_{\xi}+\text{higher order term}.$
Notice the kernel $(s-\tau)^{H-1/2}$ is unbounded when $20<H<1/$ 2 . In this case, disregarding the higher order term,
$\int_{0}^{t_{j}}$ $\left(f(s,X_{\tau^{N}(s)}^{N},Y_{s})-f(\tau^{N}(s),X_{\tau^{N}(s)}^{N},Y_{\tau^{N}(s)})\right)\;\mathrm{d}s$
$\qquad\sim\int_{0}^{t_{j}}\int_{\tau^{N}(s)}^{s}(s-\tau^{N}(s))^{H-1/2}\;\mathrm{d}W_{\xi}\;\mathrm{d}s\sim(\Delta t_{N})^{H+1/2}\int_{0}^{t_{j}}\;\mathrm{d}W_{\xi}.$
which yields a strong convergence of order $H+1/2,$ better than the previously known rate of $H.$ For $1/2\leq H<1,$ there are other terms of order 1, so that we obtain order $\min\{H+1/2,1\},$ for $10<H<$ 1 .
@@ 2. Pathwise solutions and main assumptions
For the notion and main results on pathwise solutions for RODEs, we refer the reader to [18, Section 2.1] and [30, Section 3.3]. We start with a fundamental set of conditions that imply the existence and uniqueness of pathwise solutions of the RODE Equation 1.1 in the sense of Carathéodory:
@@
We consider a function $f=f(t,x,y)$ defined on $I\times\mathbb{R}^{d}\times\mathbb{R}^{k}$ and with values in $\mathbb{R}^{d},$ and an $\mathbb{R}^{k}$-valued stochastic process $\{Y_{t}\}_{t\geq 0},$ where $I=[0,T]$, $T>0$, and $d,k\in\mathbb{N}.$ We make the following standing assumptions:
@@
    The function $f=f(t,x,y)$ is continous on $(t,x,y)\in I\times\mathbb{R}^{d}\times\mathbb{R}^{k}.$
@@
    The mapping $x\mapsto f(t,x,y)$ is globally Lipschitz continuous on $x$, uniformly on $t\in I$ and locally on $y\in\mathbb{R}^{k},$ i.e.
@@
    The stochastic vector process $\{Y_{t}\}_{t\geq 0}$ is an adapted semi-martingale on a right-continuous, filtered, and complete probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P}),$ i.e. where $\mathcal{F}_{0}$ contains all the null sets of probability distribution $\mathbb{P},$ and with a right-continuous filtration $\mathcal{F}_{t}=\cap_{s>t}\mathcal{F}_{s}$ in the $\sigma$-algebra $\mathcal{F}.$
@@
has a unique solution $t\mapsto X_{t}(\omega)$, for the realizations $X_{0}=X_{0}(\omega)$ of the initial condition and $t\mapsto Y_{t}(\omega)$ of the noise process (see [12, Theorem 1.1]). Each pathwise solution of Equation 2.2 is absolutely continuous. Moreover, the mapping $(t,\omega)\mapsto X_{t}(\omega)$ is (jointly) measurable (see [18, Section 2.1.2] and also [1, Lemma 4.51]) and, hence, give rise to a well-defined stochastic process $\{X_{t}\}_{t\in I}.$
@@
Since $\{Y_{t}\}_{t\geq 0}$ is a semi-martingale, the sample paths $t\mapsto Y_{t}(\omega)$ are almost surely continuous except for a set of measure zero (see Section 7). Then, since $f=f(t,x,y)$ is continuous, the integrand $s\mapsto f(s,X_{s}(\omega),Y_{s}(\omega))$ is (Riemann) integrable, and each pathwise solution is absolutely continuous.
@@
Since the paths of a semi-martingale are càdlàg (see the definition in Section 4), the noise process is in fact progressively measurable with respect to the filtration $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P}).$ In turn, the solution process $\{X_{t}\}_{t\in I}$ is adapted, hence also progressively measurable, to the filtration $(\Omega,\mathcal{F},(\tilde{\mathcal{F}}_{t})_{t\geq 0},\mathbb{P}),$ where each ~$\tilde{\mathcal{F}}_{t}$ is the smallest $\sigma$-algebra generated by the union of $\mathcal{F}_{t}$ and the smallest $\sigma$-algebra that makes the initial condition $X_{0}$ measurable.
Adding and subtracting $f(s,0,Y_{s})$ and using assumption Item ii yield the pathwise bound
$\|X_{t}\|\leq\|X_{0}\|+\int_{0}^{t}\left(\|f(s,0,Y_{s})\|+L(\|Y_{t}\|)\|X_{t}\|\right)\;\mathrm{d}s,$
almost surely. Using Grownwall’s inequality, we obtain the following result.
@@
This is a useful pathwise estimate on its own and is useful for pathwise convergence. For the strong convergence, however, we need further assumptions. We assume that $X_{0}$ has finite expectation, that $\|f(t,0,Y_{t})\|$ has integrable expectation, and that $L$ is bounded, i.e.
@@
Under the 2.1, assume that Equation 2.4, Equation 2.5 and Equation 2.6, hold. Then
@@
In a number of cases, depending on the structure of the equation, we might not need all the previous assumptions. Indeed, when some bounded, positively invariant region exists and is of interest, we may truncate the nonlinear term to achieve the desired global conditions for the equation with the truncated term, but which coincides with the original equation in the region of interest. In other situations, higher order moments of the noise and/or of the initial condition may compensate the lack of some of these assumptions and still yield bounds for the expectation of the norm of the solution. We leave these cases to be handled in the applications.

### v8

@@ Abstract.
It is well known that the Euler method for a random ordinary differential equation $\mathrm{d}X_{t}/\mathrm{d}t=f(t,X_{t},Y_{t})$ driven by a stochastic process $\{Y_{t}\}_{t\in I}$, on a time interval $I,$ with $\theta$-Hölder sample paths is of strong order $\theta$ with respect to the time step, provided $f=f(t,x,y)$ is sufficiently regular and with suitable bounds. This order is known to increase to $1$ in some special cases. Here, it is proved that, in many more typical cases, further structures on the noise can be exploited so that the strong convergence is of order 1. In fact, we prove so for any semi-martingale noise. This includes Itô diffusion processes, point-process noises, transport-type processes with sample paths of bounded variation, and time-changed Brownian motion. The result follows from estimating the global error as an iterated integral over both large and small mesh scales, and by switching the order of integration to move the critical regularity to the large scale. The work is complemented with numerical simulations showing the optimality of the strong order 1 convergence in those cases, and with an example with fractional Brownian motion noise with Hurst parameter $0<H<1/2,$ which is not a semi-martingale and for which the order of convergence is $H+1/2$, hence lower than the attained order 1 in the semi-martingale case, but still higher than the order $H$ of convergence expected from previous works.
@@ 1. Introduction
$\begin{cases}\displaystyle\frac{\mathrm{d}X_{t}}{\mathrm{d}t}=f(t,X_{t},Y_{t}),\\ \left.X_{t}\right|_{t=0}=X_{0},\end{cases}$ (1.1)
for $t$ on a time interval which we take to be $I=[0,T]$, with $T>0$, where the initial condition $X_{0}$ is a given random variable, and the noise $\{Y_{t}\}_{t\in I}$ is a given stochastic process, both defined on a probability space $(\Omega,\mathcal{F},\mathbb{P}).$ The function $f=f(t,x,y)$ is such that the differential equation in Equation 1.1 can be either a scalar or a system of equations and the noise can also be either a scalar or a vector-valued process.
Random ODEs appear in applications where the parameters either vary randomly or are not fully known. As such, they have been playing an important role in modeling practical problems, including uncertainty quantification (see e.g. [3, 10, 17, 20, 36, 39, 43, 44]). They are less known than the more popular Itô-diffusion stochastic differential equations (SDE), but, depending on the situation, they may actually lead to more realistic models.
@@
An approximation $\{X_{t_{j}}^{N}\}_{j=0,1,\ldots,N}$ is said to converge to $\{X_{t}\}_{t\in I}$ with strong order $\theta>0$ when there exists a constant $C\geq 0$ such that
@@
where $\mathbb{E}[\cdot]$ indicates the expectation of a random variable on the underlying probability space, and $\|\cdot\|$ is a suitable norm in the appropriate phase space.
@@
Under certain regularity conditions on $f$ and when the noise $\{Y_{t}\}_{t\in I}$ has $\theta$-Hölder continuous sample paths, the result in [48, Theorem 3] proves the order of convergence of the Euler scheme, in the mean square sense, to be of order $\theta$ with respect to the time step. This implies the strong convergence Equation 1.3 with the same order $\theta$. In the case of fractional Brownian motion noise with Hurst parameter $H$, [48, Theorem 2] proves the mean square convergence to be of order $H$, for any $0<H\leq 1.$
In some special cases, the order of convergence of the Euler method is known to be higher than the Hölder exponent of the noise. For example, for Itô diffusion noises, the RODE can be seen as part of a system of SDEs, with the Euler method for the RODE component coinciding with the Milstein method for the SDE, which is known to be of strong order 1 (see [48, Section 5], in particular [48, Section 5.2, Example 12 and Remark 13]).
Figure 1. (a) Estimated order of convergence of the strong error of the Euler method for the linear system Equation 1.4 as it depends on the noise, starting with all the noises combined (all) and continuing with the scalar equation, each with a different type of noise process, namely Wiener (W), Ornstein-Uhlenbeck (OU), geometric Brownian motion (gBm), a homogeneous linear Itô diffusion (hlp), compound Poisson process (cP), step point process (sP), Hawkes process (H), transport process (T), and also a fractional Brownian motion with Hurst parameter $1/2<H<1.$ (fBm); (b) Sample paths of all the noises used in the linear system Equation 1.4.
In many more classical examples, though, the actual order of convergence seems to be higher than what the theory predicts. Our aim here is to show that, indeed, it is possible to exploit further properties of the noise that yield a strong order 1 convergence for essentially any noise that is a semi-martingale (see Theorem 7.1 for the precise statement). This includes noises such as point processes, transport processes, Itô processes, and time-changed Brownian motions, some of which are not even continuous and for which the previous theories did not apply.
This strong order 1 convergence is not only a major improvement from previous results, but it is also remarkable that it follows with a unified proof for all semi-martingale noises. It also has potential extensions to pathwise convergence for RODEs in the case of non-globally Lipschitz functions, at least when driven by finite-variation process. The result should also extend to random partial differential equations. It is not clear, though, whether this technique could also yield improvements in the order of convergence for higher order methods.
@@
Figure 3. (a) Estimated order of convergence of the strong error of the Euler method for the toggle-switch model Equation 1.6; (b) Euler approximations of the component $\{X_{t}\}_{t\in I}$ of a sample path.
We numerically verify, in a series of examples, that the strong order 1 convergence is in fact optimal. Indeed, this work was motived by the fact that the observed convergence in these and many other examples was higher than what the theory predicted. The examples are presented in more details in Section 8 and in Appendix A and in the github repository [30]. Below is a summary of these examples:
@@
    where $\{Y_{t}\}_{t\in I}$ is either (a) a univariate process with one of many types of semi-martingale processes or (b) a vector-valued process combining all these noises as independent components.
@@
    A logistic model of population dynamics with random coefficients, loosely inspired by [20, Section 15.2], and where the specific growth is the sine of a geometric Brownian motion process $\{G_{t}\}_{t\in I}$ and with an extra point-process random term $\{H_{t}\}_{t\in I}$ representing harvest:
@@
    A toggle-switch model of gene expression (similar to [3, Section 7.8], originated from [47], see also [44]) driven by a combination of a compound Poisson point process $\{A_{t}\}_{t\in I}$ and an Itô diffusion process $\{B_{t}\}_{t\in I},$
@@
    for the surplus $U_{t}=X_{t}+C_{t}+O_{t}$ of an insurance company, inspired by [18] and [11], where $C_{t}=\sum_{i=1}^{N_{t}}C_{i}$ defines a compound Poisson process modeling the insurance claims, $\{R_{t}\}_{t\in I}$ is a geometric Brownian motion process modeling the interest rate, and $\{O_{t}\}_{t\in I}$ is an Ornstein-Uhlenbeck process, used to transform an additive SDE into a RODE.
@@ A Fisher-KPP partial differential equation
    as inspired by the works [43] and [17] (see also [16] and [33]), and where the noise $\{Y_{t}\}_{t\in I}$ is an Ornstein-Uhlenbeck process modulated by a decaying self-exciting Hawkes point process.
@@
    where $\{B^{H}_{t}\}_{t\in I}$ is a fractional Brownian motion (fBm) with Hurst parameter $0<H<1$, which is a Wiener process when $H=1/2$ and which is not a semi-martingale for $H\neq 1/2.$
Figure 4. (a) Estimated order of convergence of the strong error of the Euler method for the insurance risk model Equation 1.8; (b) Sample paths of the surplus $\{U_{t}\}_{t\in I},$ the Ornstein-Uhlenbeck process $\{O_{t}\}_{t\in I},$ the interest rate $\{I_{t}\}_{t\in I},$ and the insurance claims $\{C_{t}\}_{t\in I}.$
Figure 5. (a) Estimated order of convergence of the strong error of the Euler method for the Fisher-KPP model Equation 1.9-Equation 1.10; Sample paths of (b) the Ornstein-Uhlenbeck and Hawkes processes, and of (c) the noise obtained as the product of these processes.
In all but the last example, the numerically obtained convergence was found to be of strong order approximately 1, while previous results would either not be applicable or yield a much lower order of convergence (see Figures 1, 2, 3, 4 and 5).
@@
Figure 6. Estimated order $p$ of strong convergence of the Euler method for Equation 1.11 at different values of the Hurst parameter $H$ (scattered plot) along with the theoretical value $p=\min\{H+1/2,1\}$ proved here (dashed line), and the previously known theoretical value $p=H$ (dotted).
@@
Details about the numerical setup for estimating the order of strong convergence are given in Section 8. Details concerning the first example Equation 1.4 are given in Section 8.2. A more detailed description of the numerical setup and the other examples can be found in the open-access github repository [30].
Concerning the proof of strong order 1 convergence, the starting point is not to bound the local error and, instead, work with an explicit formula for the global error (see Lemma 5.1), similarly to the way it is done for approximations of stochastic differential equations. In this way, we obtain the bound (see Equation 6.4)
@@
The only problematic, noise-sensitive term in Equation 1.12 is the last one. The classical analysis is to use an assumed $\theta$-Hölder regularity of the noise sample paths and bound the local strong error by $C\Delta t_{N}^{\theta}.$ Instead, we look at the global noise error and show that the steps of the process given by $F_{t}=f(t,X_{\tau^{N}(t)}^{N},Y_{t})$ can be controlled in a suitable global way. In order to give the main idea, let us consider a scalar equation with a scalar noise and assume that the sample paths of $\{F_{t}\}_{t\in I}$ satisfy
@@
For a semi-martingale noise, the integral above can be a combination of a Lebesgue-Stieltjes integral, a stochastic integral, and an integral with respect to a jump measure, but we leave it like that just for the sake of explaining the main idea. With that, we bound the global error term using the Fubini Theorem, taking into consideration that, despite the inner integral ranging from $\xi$ to the “future” time $\tau^{N}(\xi)+\Delta t_{N}$, the integrand is still non-anticipative, hence
@@
In the fractional Brownian motion case, which is not a semi-martingale, the argument above does not apply. Instead, we essentially have (see Equation A.10 and Equation A.12), in the case $0<H<1/2,$
$F_{s}-F_{\tau}=\int_{\tau}^{s}g(\xi)(s-\tau)^{H-1/2}\;\mathrm{d}W_{\xi}+\text{higher order term},$
for a suitable function $g=g(\xi)$. Notice the kernel $(s-\tau)^{H-1/2}$ is unbounded when $0<H<1/2.$ In this case, disregarding the higher order term,
$\int_{0}^{t_{j}}\left(f(s,X_{\tau^{N}(s)}^{N},Y_{s})-f(\tau^{N}(s),X_{\tau^{N}(s)}^{N},Y_{\tau^{N}(s)})\right)\;\mathrm{d}s\\ =\int_{0}^{t_{j}}\int_{\tau^{N}(s)}^{s}g(\xi)(s-\tau^{N}(s))^{H-1/2}\;\mathrm{d}W_{\xi}\;\mathrm{d}s=(\Delta t_{N})^{H+1/2}\int_{0}^{t_{j}}g(\xi)\;\mathrm{d}W_{\xi},$
which yields a strong convergence of order $H+1/2,$ better than the previously known rate of $H.$ For $1/2\leq H<1,$ there are other terms of order 1, so that we obtain order $\min\{H+1/2,1\},$ for $0<H<1.$
@@ 2. Pathwise solutions and main assumptions
For the notion and main results on pathwise solutions for RODEs, we refer the reader to [20, Section 2.1] and [39, Section 3.3]. We start with a fundamental set of conditions that imply the existence and uniqueness of pathwise solutions of the RODE Equation 1.1 in the sense of Carathéodory. In what follows, $\|\cdot\|$ denotes de Euclidian norm for vectors and the corresponding operator norm for matrices.
@@
Let $d,k\in\mathbb{N}$ and $I=[0,T]$, where $T>0.$ We consider a function $f=f(t,x,y)$ defined on $I\times\mathbb{R}^{d}\times\mathbb{R}^{k}$ and with values in $\mathbb{R}^{d},$ a initial condition $X_{0}$, and an $\mathbb{R}^{k}$-valued stochastic process $\{Y_{t}\}_{t\in I}.$ We make the following standing assumptions:
@@
    The function $f=f(t,x,y)$ is continuous in $(t,x,y)\in I\times\mathbb{R}^{d}\times\mathbb{R}^{k}.$
@@
    The mapping $x\mapsto f(t,x,y)$ is globally Lipschitz continuous in $x$, uniformly in $t\in I$ and locally in $y\in\mathbb{R}^{k},$ i.e.
@@
    The stochastic vector process $\{Y_{t}\}_{t\in I}$ is an adapted semi-martingale on a right-continuous, filtered, and complete probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P}),$ i.e. where $\mathcal{F}_{0}$ contains all the null sets of probability distribution $\mathbb{P},$ and with a right-continuous filtration $\mathcal{F}_{t}=\cap_{s>t}\mathcal{F}_{s}$ in the $\sigma$-algebra $\mathcal{F}.$
@@
has a unique solution $t\mapsto X_{t}(\omega)$, for the realizations $X_{0}=X_{0}(\omega)$ of the initial condition and $t\mapsto Y_{t}(\omega)$ of the noise process (see [13, Theorem 1.1]). Each pathwise solution of Equation 2.2 is absolutely continuous. Moreover, the mapping $(t,\omega)\mapsto X_{t}(\omega)$ is (jointly) measurable (see [20, Section 2.1.2] and also [2, Lemma 4.51]) and, hence, gives rise to a well-defined stochastic process $\{X_{t}\}_{t\in I}.$ Equation (2.2) is the integral form of equation (1.1), so that $\{X_{t}\}_{t\in I}$ is a solution of (1.1) in the pathwise sense.
@@
Since $\{Y_{t}\}_{t\in I}$ is a semi-martingale, the sample paths $t\mapsto Y_{t}(\omega)$ are almost surely continuous except for a set of measure zero (see Section 7). Then, since $f=f(t,x,y)$ is continuous, the integrand $s\mapsto f(s,X_{s}(\omega),Y_{s}(\omega))$ is (Riemann) integrable, and each pathwise solution is absolutely continuous.
@@
Since the paths of a semi-martingale are càdlàg (see the definition in Section 4), the noise process is in fact progressively measurable with respect to the filtration $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P}).$ In turn, the solution process $\{X_{t}\}_{t\in I}$ is adapted and hence also progressively measurable to the filtration $(\Omega,\mathcal{F},(\tilde{\mathcal{F}}_{t})_{t\geq 0},\mathbb{P}),$ where each $\tilde{\mathcal{F}}_{t}$ is the smallest $\sigma$-algebra generated by the union of $\mathcal{F}_{t}$ and the smallest $\sigma$-algebra that makes the initial condition $X_{0}$ measurable.
Adding and subtracting $f(s,0,Y_{s})$ and using assumption (ii) of the 2.1 yield the pathwise bound
$\|X_{t}\|\leq\|X_{0}\|+\int_{0}^{t}\left(\|f(s,0,Y_{s})\|+L(\|Y_{s}\|)\|X_{s}\|\right)\;\mathrm{d}s,$
almost surely. Using Gronwall’s inequality, we obtain the following result.
@@
This is a useful pathwise bound on its own and is particularly useful for proving pathwise convergence. For the strong convergence, however, we need further assumptions. We assume that $X_{0}$ has finite expectation, that $\|f(t,0,Y_{t})\|$ has integrable expectation, and that $L$ is bounded, i.e.
@@
Under the 2.1, assume that Equation 2.4, Equation 2.5 and Equation 2.6 hold. Then
@@
In a number of cases, depending on the structure of the equation, we might not need all the previous assumptions, especially the restrictive assumption (2.6), corresponding to a global Lipschitz condition in $x,$ uniformly in $t$ and $y$. Indeed, when some bounded, positively invariant region exists and is of interest, we may truncate the nonlinear term to achieve the desired global conditions. Within the invariant region, the equation with the truncated term coincides with the original equation, and so do their Euler approximations, for a sufficiently refined mesh. In other situations, higher order moments of the noise and/or of the initial condition may compensate the lack of some of these assumptions and still yield bounds for the expectation of the norm of the solution. In the examples described in the Introduction, the linear models (iv) and (vii) are the only ones that satisfy the global condition (2.6). All the other examples fall within the framework of having a bounded, positively invariant region. An example in which neither applies but bounds on higher order moments compensate is, for example, a linear equation with an unbounded multiplicative noise, which can be found in the github repository [30]. Finally, it is worth mentioning that the global condition (2.6) is the classical assumption used for the strong convergence of approximatinos of stochastic differential equations [4, 19, 20, 23, 26, 31].